# A Neural-Guided Variational Quantum Algorithm for Efficient Sign Structure Learning in Hybrid Architectures

### v1

@@ Results
For sufficiently large $L$, $\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left(f^{2}(s)\right)<1$. By the uniform convergence of the series,
$\text{Var}(g^{-1}(s,L))=\sum_{i}\mu_{i}\text{Var}^{i}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\left[f^{2}(s)\right],$ (4)
with $\mu_{i}>0$. This variance asymptotically approaches zero as $L\to\infty$. In our numerical experiment, as illustrated in Figures 3 and 4, the coefficient of variation of energy $\text{CV}\propto\sqrt{Var(g^{-1}(s,L))}$ decays with $L$, suggesting that the algorithm has better noise immunity under actual sampling constraints as $L$ increase. In summary, we eliminate a source of numerical instability (which subsequently affects the convergence rate) of the original VQNHE.
@@
To rigorously validate the accuracy and efficiency of our proposed algorithm for the ground-state preparation of a given Hamiltonian, we conducted systematic evaluations across four canonical manybody problems: the 1D J1-J2 model, the 2D Heisenberg model, the Transverse Field Ising Model (TFIM), and the Classical Ising model. These models were chosen to represent diverse physical systems with varying degrees of sign problem severity and ground-state complexity. We employed three methodologies, sVQNHE, VQE, and NNs, to benchmark their performance. Additionally, we integrated the encoding scheme proposed in Ref. [36] to demonstrate the superiority of sVQNHE over conventional VQAs (such as QAOA and VQE) in solving the MaxCut problem, a prototypical combinatorial optimization task.
Our experiments included a 6-qubit 1D J1-J2 model with $J1/J2=0.6$, a configuration known to exhibit a highly non-trivial ground-state phase structure [45]. This model serves as a representative case for systems with intricate sign signatures. We also tested a 9-qubit 2D Heisenberg model with periodic boundary conditions, which is known to incur severe sign-related challenges for numerical methods[46]. For comparison, we analyzed a 9-qubit 1D TFIM (solvable via Jordan-Wigner transform) and two 1D Ising models (6 and 12 qubits). Both TFIM and Ising models we chose are Sign-problem-free systems. Nevertheless, they still present non-trivial challenges in different ways. For example, they are relevant to the qubit-formulated combinatorial optimization problems. We evaluated our algorithm under four scenarios to assess its theoretical advantages and practical robustness:
  • Ideal Simulation: Quantified intrinsic accuracy, variability, and scalability under ideal conditions.
  • Real-World Simulation: Assessed robustness under realistic quantum conditions, combining depolarization-based gate errors (two-qubit: $0.005$; single-qubit: $0.0013$) with bitstring-sampling-based estimation of observables and gradients.
  • Scalability Under Finite Sampling: Demonstrated that sVQNHE’s advantage over existing methods[36] can be sustained with system size.
We emphasize that the results of all the so-called "non-ideal" experiments are based on the original sampling data (of bitstrings) from noisy quantum circuits, without implementing any quantum error mitigation procedures. As sVQNHE possesses an inherent and robust error mitigating capability [47] due to the fact that the hybrid quantum-neural ansatz can be directly tuned with the classical NN module during the noisy optimization process.
1. Ideal Simulation
We conducted a systematic evaluation of the performance of sVQNHE and other methods (VQNHE, VQE, NN, QAOA) focusing on accuracy, variability, and measurement efficiency. All the calculations were done in statevector-type exact numerical computations. For consistency, sVQNHE and VQNHE utilized the same non-negative real neural network component (MLP), while the NN method employed a real MLP with identical parameters and structure. The quantum components of sVQNHE and VQNHE were our single-layer sign ansatz and a hardware-efficient ansatz with a similar number of parameters, respectively. In Table 2, these are denoted as sign-NN and hea-NN, respectively. The VQE method used a hardware-efficient ansatz with varying layers, such as HEA2 for two layers. Similarly, the QAOA method was tested with different layers, for instance, QAOA10 for ten layers. All methods initialized quantum circuit parameters using independent uniform distributions $U(-2\pi,2\pi)$, optimized for the same number of steps, and averaged the estimated energy from 20 random initial points. The baseline algorithm in Table 2 is NN.
Simulating Frustrated Many-Body Systems
For the 1D J1-J2 model, which features a non-trivial ground state, the sVQNHE method (sign-NN) demonstrated significant improvements over the NN method, the VQE method (HEA2), and the VQNHE method (hea-NN). Specifically, Table 2 shows that after 2000 optimization steps, the sign-NN method reduced the mean absolute error (MAE) by $98.9\%$ and the variance by $99.6\%$ for the 6-qubit J1-J2 model compared to the NN method. Additionally, Figure 2 illustrates the scaling results (from 6 qubits to 12 qubits) of sign-NN and NN with the same number of optimization steps to achieve the minimum error. The advantage of sign-NN increases with the number of qubits. In contrast, while hea-NN also reduced the MAE by $75.1\%$ compared to NN, its variance was larger than that of NN (variance: hea-NN > NN > sign-NN). Furthermore, due to the non-commutative structure of hea, the measurement amount for sign-NN was significantly smaller than that for hea-NN under similar parameter amounts. These three points—improved accuracy, reduced variance, and reduced measurement amount—demonstrate the superiority of sVQNHE over the original VQNHE algorithm. Regarding convergence speed, sign-NN significantly outperformed HEA2. To achieve the same error threshold, HEA2 required 12.6 times more optimization steps on average than sign-NN. In the most efficient cases, HEA2 needed 1676 steps to reach the same accuracy, while sign-NN only required 89 steps. Additionally, the probability of the sign-NN successfully reaching the target accuracy was 12 times higher than that of HEA2, highlighting its excellent reliability and efficiency (for details, please refer to the supplementary material). Similar results were observed in the Heisenberg model, further verifying the capability of sVQNHE to handle systems affected by the sign problem.
Model Method $R_{MAE}$ $R_{Var}$
6-qubit 1D J1-J2 sign-NN -0.989 -0.996
hea-NN -0.751 1.955
HEA2 -0.916 -0.132
9-qubit 2D Heisenberg HEA1 -0.331 0.554
hea-NN -0.119 33.793
sign-NN -0.872 -0.832
9-qubit 1D TFIM HEA1 -0.290 0.019
sign-NN -0.640 -0.728
9-qubit 1D Ising HEA1 0.129 0.909
QAOA5 20.674 15.583
QAOA10 1.205 11.835
sign-NN -0.416 -0.455
12-qubit 1D Ising HEA1 6.822 10.002
QAOA10 77.012 80.404
sign-NN -0.811 -0.979
@@ 1. Ideal Simulation
Table 2: Performance Analysis of sVQNHE Algorithm. Here, the range of $R_{MAE}$ and $R_{Var}$ is $(-1,+\infty)$. The closer to $-1$, the better (see Equations (12) and (13) in the Method section for details). The baseline algorithm NN model is a real Multi-Layer Perceptron (MLP). HEA1 and HEA2 represent VQE methods employing 1-layer and 2-layer Hardware-Efficient Ansatz, respectively. QAOA5 and QAOA10 represent the QAOA method with layers 5 and 10, respectively. The Sign Ansatz employs $R_{zz}$ gates to connect all edges of the corresponding physical model.
@@
Figure 2: Scaling test for the 1D J1-J2 model. The x-axis shows the number of qubits, while the y-axis displays the average relative error from the sVQNHE method (sign-NN, blue line) and the NN method (NN, orange line) after 20 runs of 2000 steps each. The green line illustrates the difference in average relative error between the NN and sign-NN methods as the qubit count increases. Shaded regions represent the relative error bar scaled by 0.5.
Simulating Sign-Problem-Free Many-Body Systems
Although the above results highlight the advantages of sVQNHE in sign-frustrated quantum models, the same architectural design is effective even in sign-free systems, where phase modeling is trivial for the ground state itself, but the ability to explore a broader Hilbert space (as defined by the PQC-generated phase strcutures) turns out to be beneficial for the optimization convergence in the numerical investigated presented in this subsection. More precisely, please see the last three rows of Table 2.
In our numerical experiments on the Ising model, the sign-NN approach demonstrates substantial improvements over the NN in both error reduction and variance compression. At 9 qubits, sign-NN achieves a $41.6\%$ reduction in MAE compared to NN, indicating a significant enhancement in accuracy. This advantage becomes even more pronounced at 12 qubits, where sign-NN reduces MAE by $81.1\%$ relative to NN, showcasing its superior scalability with increasing system size. Similarly, in terms of variance, sign-NN outperforms NN by reducing variance by $45.5\%$ at 9 qubits. This gap widens dramatically at 12 qubits, with sign-NN achieving a $97.9\%$ reduction in variance compared to NN, highlighting its ability to deliver consistent and reliable results as the system scales. When benchmarked against the QAOA in the 9-qubit Ising model, sign-NN maintains its performance edge. A 5-layer QAOA configuration requires a comparable number of per-iteration measurements to sign-NN but yields a higher MAE, indicating lower accuracy. Extending QAOA to 10 layers improves its accuracy, yet it still falls short of sign-NN’s performance. This trend persists and becomes more evident at 12 qubits, where sign-NN’s scalability further amplifies its advantage over QAOA. The consistency of these results extends to the TFIM. The superior performance of sign-NN can be attributed to its carefully designed ansatz, optimized with neural network guidance. Unlike QAOA, which uses $R_{x}$ gates and constrains all $Z$- and $X$-related gates within a layer to share a single parameter, the Sign Ansatz employs $R_{zz}$ and $R_{y}$ gates with independent parameters. This design, as supported by Theorem 1, generates a higher-dimensional Lie algebra ($\mathfrak{g}_{1}$) compared to QAOA’s ($\mathfrak{g}_{2}$), enabling sign-NN to explore a broader unitary space through multi-qubit $Z$-phase adjustments. The phase-focusing architecture of the Sign Ansatz, which leverages multi-qubit $Z$-gates to refine phase information, allows for efficient navigation of high-dimensional solution spaces. When combined with neural network optimization, this approach achieves high accuracy and low variance at a reduced computational cost, making sVQNHE particularly well-suited for large-scale quantum optimization problems.
2. Real-World Simulation
Next, we consider how robustly sVQNHE performs when there exists both gate errors (for e.g. depolarization channels) and uncertainty induced by the finite sampling of bitstrings instead of statevector calculations as in the previous section. We evaluated sVQNHE against VQE, adopting two distinct optimization strategies, layered-VQE (layerwise optimization) and standard-VQE (joint optimization of all parameters), using an identical ansatz circuit. sVQNHE employed non-negative real neural networks, while both VQE methods utilized the same Sign Ansatz without NN augmentation. For all VQE schemes, the resource consumption was saved by measurement based on commuting groups. All simulations capped iterations at 400.
Figure 3: Performance evaluation of sVQNHE and VQE variants under a 3-qubit 1D Heisenberg Hamiltonian. Energy (y-axis) versus iteration count (x-axis), with the ground state energy marked by a black dashed line. For sVQNHE and layered-VQE (layered optimization), blue and orange lines correspond to the first and second evolutionary layers, respectively, where each layer introduces new variational parameters. For standard-VQE (overall optimization), the lines reflect energy progression across sequential steps under fixed parameterization. For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01618$ and $cv_{2}=0.015912$, indicating a reduction in variability in the second layer.
@@ 2. Real-World Simulation
Figure 4: Energy convergence of sVQNHE and VQE variants under a noisy 3-qubit 1D Heisenberg Hamiltonian. Iteration count (x-axis) versus energy (y-axis). The solid black dashed line marks the ground state energy value. For layered optimization (sVQNHE, layered-VQE), blue and orange lines denote first- and second-layer parameter updates, respectively; standard-VQE (overall optimization) reflects sequential parameter refinement. For sVQNHE, coefficients of variation after stabilization (final 50 iterations per layer) are $cv_{1}=0.01534$ and $cv_{2}=0.01487$, indicating a reduction in variability in the second layer.
As shown in Figures 3 and 4, sVQNHE consistently achieved faster convergence than VQE, both with and without device noise. While layered optimization inherently reduces computational resource demands, mitigates noise sensitivity via shallower circuits, and often accelerates convergence, it is not applicable to the general VQE framework due to premature convergence to local optimal solutions. Specifically, the 1-layer Sign Ansatz fails as a viable initial parameterization for VQE’s layered optimization, as evidenced by the divergent performance of layer-VQE and standard VQE in our tests. However, integrating NNs into the Sign Ansatz (as in sVQNHE) circumvents this limitation, enabling effective layered optimization by delegating partial modeling duty to the neural network. Post-convergence analysis of sVQNHE revealed that the coefficient of variation (CV) for the second optimization layer remained lower than that of the first layer across all noise conditions. This validates the efficacy of our gradual transfer mechanism, which systematically refines solutions while maintaining stability under noise.
3. Scaling Under Sampling Errors
Finally, we investigate whether the observed advantage of sVQNHE can be sustained as the system size increases under the real-world setting of extracting PQC state by sampling bitstrings on quantum circuits. An example is provided in Figure 5. We evaluate the solution quality and measurement cost of the MaxCut problem on Erdős–Rényi random graphs with an edge connection probability of $0.3$ as the graph size increases. To establish this advantage, we compare sVQNHE against the Sign Ansatz VQE (sign-VQE) and the recently proposed brickwork-VQE [36], all evaluated under the same polynomial encoding model. This model encodes the $m$ binary variables of the MaxCut problem into the Pauli correlations across $k$ qubits, enabling the representation of $m=O(n^{k})$ variables with only $n$ qubits, where $k>1$ is an integer. Unlike the direct encoding schemes like QAOA, which require $m$ qubits—one per variable—this approach achieves polynomial compression in the qubit count. The desired solution of the combinatorial problem is now encoded in terms of a non-classical Hamiltonian. The compression ratio depends on the polynomial order $k$, allowing us to tackle larger MaxCut instances with fewer qubits. All methods are evaluated under a unified encoding framework with $k=2$ to ensure fairness.
To evaluate solution quality across varying scales, we conducted experiments with 1,200 measurement shots, averaging results over multiple trials. As shown in Figure 5(Left), scaling sVQNHE parameters—through enhancements in the NN component or other optimizations—consistently widens its performance gap over competing methods. Specifically, sVQNHE, which employs NN-guided hierarchical optimization and NN post-processing, significantly outperforms sign-VQE and brickwork-VQE. Sign-VQE adopts the final PQC structure from sVQNHE (a two-layer sign ansatz with two diagonal and two non-diagonal layers) but omits the NN-guided optimization and post-processing used in sVQNHE. Brickwork-VQE, in contrast, utilizes a VQE with a brickwork ansatz, as detailed in [36]. Figure 5(Left) reveals that, without NN guidance and post-processing (sign-VQE vs. brickwork-VQE), the sign ansatz shows no clear advantage over the brickwork ansatz, with comparable solution quality and similar measurement costs. However, sVQNHE’s hybrid approach - integrating hierarchical optimization guided by NN and NN post-processing to Sign Ansatz - yields a marked improvement. For instance, the performance gap between sVQNHE and baseline methods increases by approximately $6\%\approx(0.96-0.90)/0.90$ for 9-vertex graphs and $19\%\approx(1-0.84)/0.84$ for 45-vertex graphs when using parameter set 2. Additionally, Figure 5(Right) shows that sVQNHE significantly reduces single-step measurement consumption compared to both sign-VQE and brickwork-VQE. With parameter set 2, it achieves $86\%\approx(72-10)/72$ and $85\%\approx(306-46)/306$ improvements in quantum resource efficiency for 9- and 45-vertex systems, respectively. Additionally, in the Supplementary Information, we provide a preliminary comparison of resource consumption for larger instances, further highlighting sVQNHE’s advantages in scalability and efficiency under large-scale conditions. This underscores its potential for large-scale quantum applications.

### v2

@@ Results
For sufficiently large $L$, the variance of $f^{2}(s)$ becomes small ($\text{Var}(f^{2}(s))<1$), which implies that $g(s,L)$ concentrates tightly around its mean $\mathbb{E}[g(s,L)]=\mathbb{E}\left[f^{2}(s)\right]$.
Under this concentration condition, the variance of the normalization factor $g^{-1}(s,L)$ can be accurately approximated using the first-order delta method (also known as the linearization method):
$\text{Var}\bigl(g^{-1}(s,L)\bigr)\approx\frac{1}{\mathbb{E}^{4}[g(s,L)]}\text{Var}(g(s,L))=\frac{1}{\mathbb{E}^{4}[g(s,L)]}\text{Var}_{s\sim|\langle s|\psi_{L}\rangle|^{2}}\bigl[f^{2}(s)\bigr].$ (4)
[…]
In our numerical experiments, as illustrated in Figures 4(a) and (b), the coefficient of variation of the energy estimate $\text{CV}(\hat{E})\propto\sqrt{\text{Var}(g^{-1}(s,L))}$ decays with increasing $L$. This confirms that sVQNHE exhibits improved noise immunity and robustness under realistic finite-sampling constraints as the quantum circuit depth grows. In summary, we mitigate a source of numerical instability of the original VQNHE.
@@
To systematically evaluate the proposed sVQNHE framework, we design a 4-stage benchmark suite that progressively demonstrates its key advantages: (1) its ability to overcome optimization challenges arising from complex phase structures in strongly frustrated quantum many-body systems, (2) its general performance improvement even in systems with trivial phase structures due to enhanced optimization landscapes and measurement efficiency, (3) its robustness and noise mitigation capabilities under real NISQ conditions (without any error mitigation), and (4) its potential to deliver near-term quantum utility in large-scale combinatorial optimization using a qubit-efficient encoding.
All quantum simulations are performed without any form of quantum error mitigation. Noisy results are obtained directly from shot-based sampling of noisy quantum circuits. Unless otherwise stated, energy values are reported as relative errors with respect to the exact ground-state energy, and all results are averaged over multiple random initializations.
1. Frustrated Quantum Many-Body Systems
The experimental application of the sVQNHE framework to frustrated quantum spin systems serves as a critical validation of its advantages. These systems, characterized by highly complex ground-state phase structures and strong correlations, pose significant challenges to conventional variational quantum algorithms. Because it requires deep circuits to optimize both amplitude and phase components simultaneously, they often encounter optimization challenges, such as getting stuck in barren plateaus or local minima [mcclean2018barren, cao2025exploiting]. By explicitly separating these tasks, delegating phase learning to a shallow quantum module while modeling the amplitude with classical neural networks, sVQNHE effectively mitigates the complex phase structures problem. The resulting faster and more stable convergence demonstrates that sVQNHE successfully alleviates key optimization dilemmas inherent to traditional VQEs, thereby achieving superior performance in strongly correlated systems compared to purely classical or purely quantum approaches and highlighting the framework’s potential as an efficient hybrid strategy for tackling complex quantum many-body problems.
We focus on two paradigmatic models with highly non-trivial ground-state phase diagrams and strong sign oscillations in the computational basis to test:
    the 1D J1–J2 chain with $J_{1}/J_{2}=0.6$ (in the Majumdar–Ghosh regime near the dimerized phase, known for its challenging phase structure [viteritti2022accuracy]),
    the 2D square-lattice Heisenberg antiferromagnet with periodic boundary conditions (9 qubits, exhibiting severe phase complexity in numerical methods [stoudenmire2012studying]).
We compare sVQNHE (sign-VQNHE) against four strong baselines:
    pure classical neural network (NN) with real-valued MLP,
    hardware-efficient VQE (HEA) with 1 or 2 layers,
    original VQNHE using the same hardware-efficient quantum circuit and real-valued NN (denoted hea-NN),
    QAOA with up to 10 layers.
All methods are trained with the same number of optimization steps (2000) and evaluated over 20 random initializations. Performance is quantified using relative mean absolute error ($R_{\text{MAE}}$) and relative variance ($R_{\text{Var}}$) with respect to the pure NN baseline (see Methods for definitions; values closer to -1 indicate better performance). Results are summarized in Table 1.
Metric 6q J1-J2 9q Heisenberg 9q TFIM 9q Ising 12q Ising
sign-VQNHE hea-NN HEA2 HEA1 hea-NN sign-VQNHE HEA1 sign-VQNHE HEA1 QAOA5 QAOA10 sign-VQNHE HEA1 QAOA10 sign-VQNHE
$R_{\text{MAE}}$ -0.989 -0.751 -0.916 -0.331 -0.119 -0.872 -0.290 -0.640 0.129 20.674 1.205 -0.416 6.822 77.012 -0.811
$R_{\text{Var}}$ -0.996 1.955 -0.132 0.554 33.793 -0.832 0.019 -0.728 0.909 15.583 11.835 -0.455 10.002 80.404 -0.979
@@ 1. Frustrated Quantum Many-Body Systems
Table 1: Performance Analysis of sVQNHE Algorithm. Here, the range of $R_{MAE}$ and $R_{Var}$ is $(-1,+\infty)$. The closer to 1-1, the better (see Equations (LABEL:eq:R_MAE) and (LABEL:eq:R_var) in the Method section for details). The baseline algorithm NN model is a real Multi-Layer Perceptron (MLP). HEA1 and HEA2 represent VQE methods employing 1-layer and 2-layer Hardware-Efficient Ansatz, respectively. QAOA5 and QAOA10 represent the QAOA method with layers 5 and 10, respectively. The Sign Ansatz employs $R_{zz}$ gates to connect all edges of the corresponding physical model.
For the 6-qubit J1-J2 chain, sVQNHE achieves dramatic improvements. Specifically, $R_{\text{MAE}}=-0.989$ implies $98.9\%$ error reduction vs. NN, and $R_{\text{Var}}=-0.996$ implies $99.6\%$ variance reduction vs. NN. In contrast, hea-NN only reduces MAE by $75.1\%$ while increasing variance compared to NN. Conventional 2-layer HEA (HEA2) reaches -0.916 in MAE but still lags far behind sVQNHE in both accuracy and stability. Moreover, sVQNHE converges significantly faster: it requires on average $12.6\times$ fewer steps than HEA2 to reach the same error threshold, and succeeds in reaching target accuracy $12\times$ more frequently. In the most efficient cases, HEA2 needed 1676 steps to reach the same accuracy, while sVQNHE only required 89 steps (for details, please refer to the supplementary material). Furthermore, due to the non-commutative structure of hardware-efficient ansatze, the measurement overhead for sVQNHE is significantly lower than for hea-NN under similar parameter counts.
The scaling behavior is even more striking. Figure 2 shows the performance gap between sVQNHE and NN as system size increases from 6 to 12 qubits under a fixed optimization budget. The advantage of sVQNHE widens systematically with system size, demonstrating that the phase-focused architecture becomes increasingly valuable as correlation strength and phase complexity intensify.
@@
Figure 2: Scaling test for the 1D J1-J2 model. The x-axis shows the number of qubits, while the y-axis displays the average relative error from the sVQNHE method (sign-VQNHE, blue line) and the NN method (NN, orange line) after 20 runs of 2000 steps each. The green line illustrates the difference in average relative error between the NN and sign-VQNHE methods as the qubit count increases. Shaded regions represent the relative error bar scaled by 0.5.
Similar qualitative behavior is observed on the 9-qubit 2D Heisenberg model ($R_{\text{MAE}}=-0.872$, $R_{\text{Var}}=-0.832$), confirming that sVQNHE consistently resolves the trainability barriers faced by both pure classical NN (which cannot model phases) and hardware-efficient variational circuits (which entangle amplitude and phase learning) in regimes with complex phase structures.
2. Trivial Phase Structures Systems
Having demonstrated sVQNHE’s strength in systems with complex phase structures, we next ask whether the method remains advantageous in systems where the ground-state phase is trivial (i.e., real-valued or easily captured without specialized phase modeling).
We consider three representative cases without significant phase frustration:
    1D Transverse Field Ising Model (TFIM, 9 qubits, exactly solvable via Jordan–Wigner transformation),
    1D classical Ising chains (9 and 12 qubits, relevant to combinatorial optimization encodings),
    electronic ground state of the water molecule ($H_{2}O$) in STO-3G basis (10 qubits after active-space reduction, introducing electronic correlation challenges).
Table 1 shows that even without complex phases, sVQNHE systematically outperforms the pure NN baseline. Notably, the performance margin increases with system size, indicating improved scalability. When benchmarked against QAOA (5 or 10 layers), sVQNHE maintains its edge: QAOA requires more measurements per iteration and yields higher errors, even at deeper layers. The superior performance stems from sVQNHE’s phase-focusing architecture (using $R_{zz}$ and $R_{y}$ gates with independent parameters), which generates a higher-dimensional Lie algebra compared to QAOA (as per Theorem 1), enabling broader exploration of the unitary space.
For the $H_{2}O$ molecule, we compare sVQNHE against standard 3-layer hardware-efficient VQE across multiple geometries (generated with PySCF, bond length $1.2$, bond angles $90^{\circ},120^{\circ}]$). Both methods use comparable total quantum parameter counts. As shown in Figure 3, sVQNHE consistently achieves: lower normalized energy error and higher wavefunction fidelity. Across the entire geometry scan, despite using half as many quantum parameters per iteration and benefiting from fully-commuting layers (leading to dramatically lower measurement cost).
These results demonstrate that the division of the amplitude–phase task between quantum and classical resources with neural-guided optimization significantly improves trainability and solution quality even when the phase learning is not particularly challenging. The benefits arise from a smoother optimization landscape, reduced variance, and enhanced measurement efficiency, advantages that are broadly relevant across many problem classes.
Figure 3: Comparison of normalized energy and fidelity for $H_{2}O$ molecules at various bond angles (angle). The x-axis labels denote different geometric configurations (r, angle), where (r) is the bond length. The top panel (a) presents the average normalized energy $E_{\mathrm{pred}}/E_{\mathrm{exact}}$ for each group, while the bottom panel (b) shows the average fidelity between the predicted and exact wavefunctions. Blue bars represent results from the VQE method, and orange bars correspond to the sVQNHE method. Error bars indicate 0.5 standard deviation over 5 independent runs.
3. Robustness Under Noisy Conditions
To assess the practical viability of sVQNHE in the NISQ regime, we evaluate its performance under two realistic sources of imperfection: (1) finite sampling shot noise from bitstring measurements, and (2) hardware gate and readout noise (modeled via depolarizing channels). All results are obtained directly from noisy circuit sampling without applying any quantum error mitigation techniques.
We focus on the 3-qubit 1D Heisenberg Hamiltonian as a representative frustrated system. We compare sVQNHE against two VQE variants using the identical Sign Ansatz: layered optimization VQE (layerwise-VQE) and standard optimization VQE (standard-VQE). sVQNHE employs non-negative real-valued neural networks for amplitude correction, while both VQE baselines optimize only quantum parameters. All simulations are capped at 400 iterations, with resource-efficient commuting-group measurements for VQE.
Figure 4(a) shows convergence curves (energy vs. iteration) under finite sampling (ideal noise-free circuits, limited shots). sVQNHE converges faster and more stably than both VQE variants. Notably, layered optimization fails for pure VQE because the 1-layer Sign Ansatz leads to premature trapping in poor local minima. In contrast, sVQNHE successfully leverages layered growth by delegating initial amplitude modeling to the NN, enabling effective progressive refinement. Post-convergence, the coefficient of variation (CV) in the second layer is lower than in the first ($CV_{1}=0.01618\to CV_{2}=0.015912$), confirming that the gradual transfer mechanism enhances stability under sampling variance.
(a) Performance evaluation under finite sampling.
(b) Performance evaluation under noisy conditions.
@@ 3. Robustness Under Noisy Conditions
Figure 4: Comparative performance of sVQNHE and VQE variants under a 3-qubit 1D Heisenberg Hamiltonian. (a) Finite sampling (1200 shots); (b) Noisy conditions (1200 shots, two-qubit: $0.005$; single-qubit: $0.0013$). Iteration count (x-axis) versus energy (y-axis). The solid black dashed line marks the ground state energy value. For layered optimization (sVQNHE, layered-VQE), blue and orange lines denote first- and second-layer parameter updates, respectively; standard-VQE (overall optimization) reflects sequential parameter refinement. For sVQNHE, coefficients of variation after stabilization are $cv_{1}=0.01618,cv_{2}=0.015912$ (a) and $cv_{1}=0.01534,cv_{2}=0.01487$ (b), indicating reduced variability in the second layer.
Under realistic device noise (Figure 4(b)), sVQNHE again achieves faster convergence to near-ground-state energies and exhibits lower variability in the refined second layer ($CV_{1}=0.01534\to CV_{2}=0.01487$). This demonstrates that the hybrid quantum-neural architecture possesses an intrinsic error-mitigating capability. The classical NN module can be continuously tuned during noisy optimization to partially compensate for circuit-induced errors, as previously analyzed in related variational hybrid frameworks [zhang2023variational].
These noisy experiments highlight that sVQNHE not only maintains its advantages in ideal settings but also exhibits superior robustness and reliability under the combined effects of shot noise and incoherent errors.
4. Toward Near-Term Quantum Utility in Large-Scale Combinatorial Optimization
Finally, we investigate whether the sVQNHE framework can deliver meaningful quantum utility in practically relevant combinatorial optimization problems using highly compressed qubit encodings.
We consider two canonical NP-hard problems:
    MaxCut Problem (Erdős–Rényi graphs, up to 1485 vertices),
    Maximum Clique Problem (Erdős–Rényi graphs, 135 vertices).
Both problems are encoded using Pauli Correlation Encodings (PCE) [sciorilli2025towards], achieving polynomial qubit compression: $O(n^{k})$ classical variables are mapped to only $n$ qubits with $k=2$ (MaxClique) or $k=4$ (large MaxCut).
For MaxCut, under noisy sampling conditions, as shown in Figure 5 (a), sVQNHE consistently outperforms the baselines in solution quality. The performance gap widens with scale: for $m=9$, sVQNHE achieves $\sim 6\%$ improvement over baselines (e.g., 0.96 vs. 0.90), rising to $\sim 19\%$ for $m=45$ (1.00 vs. 0.84) using optimized parameter set 2 (enhanced NN depth and learning rate). Without NN components, sign-VQE and brickwork-VQE perform comparably, indicating that the sign ansatz alone offers no inherent edge over brickwork; the gains stem from sVQNHE’s hybrid integration. Figure 5 (b) highlights resource efficiency: sVQNHE reduces single-step measurement costs by $\sim 86\%$ for $m=9$ (10 vs. 72) and $\sim 85\%$ for $m=45$ (46 vs. 306) compared to baselines. This efficiency arises from NN-guided layer-wise optimization, which minimizes the number of gradient evaluations. Preliminary results in Supplementary Information extend this to larger instances ($m>45$), confirming sVQNHE’s scalability under noisy sampling.
For large-scale instances, we use state-vector simulations on graphs with $m=1485$ vertices and edge probability $0.01$, encoded via quartic PCE ($k=4$), requiring $n=12$ qubits. This setup tests the case where direct encodings are infeasible in the NISQ and early FTQC eras. We generate five random instances and compare sVQNHE against: (1) Free-Energy Machine [shen2025free] (FEM), a physics-inspired classical solver minimizing variational free energies via gradient descent; (2) D-Wave’s Simulated Annealing Sampler [dwave_ocean_sdk] (SA); and (3) a pure NN solver, using only the classical components of the corresponding sVQNHE method, without the quantum part. For each instance, as in the sampling case, we compute the cut value for all methods and normalize relative to the maximum cut across the four solvers. Averaging over the five graphs yields average relative solution qualities: sVQNHE (0.9989), FEM (0.9989), NN (0.9800), and SA (0.8066) (see Figure 5 (c)). sVQNHE and FEM deliver near-optimal solutions, with sVQNHE slightly outperforming FEM by exploiting quantum correlations for finer exploration of the rugged landscape. The pure NN lags due to underparameterization in high-dimensional spaces, while SA struggles with sparse graphs (low edge probability), as annealing heuristics favor denser structures [johnson1989optimization, myklebust2015solving, hamerly2019experimental].
For MaxClique, we evaluate on graphs with $m=135$ vertices and edge probability $p=0.5$. Encoding uses PCE with $k=2$, compressing to $10$ qubits. Methods include sVQNHE, its pure real NN component (identical classical part without quantum module), VQNHE (hardware-efficient ansatz with equivalent number of quantum parameters), greedy (degree-based iterative addition), and FEM. Non-greedy methods require post-processing (thresholding + greedy repair) to ensure valid cliques, as continuous relaxations may yield infeasible assignments. Over 3 instances, average relative solution qualities, normalized to the maximum across methods) rank as: sign-VQNHE ($1.00$), Greedy ($0.89$), NN and VQNHE ($0.81$), FEM ($0.26$) (see Figure 5 (d)). sVQNHE outperforms greedy’s typical $maxclique\approx\log_{2}135$ (though greedy requires no post-processing, its local search limits global exploration). The superior performance of sVQNHE highlights the critical value of its quantum module. Compared to the pure NN baseline (which uses the exact same classical neural network component but lacks the quantum phase-learning layers), sVQNHE delivers significantly higher solution quality, demonstrating that the quantum part (diagonal gates $W_{i}$ for phase structure learning) actively contributes meaningful computational power rather than merely consuming resources. This advantage is further validated against VQNHE (with equivalent parameter count): sVQNHE’s innovative co-learning procedure of phase and amplitude components with adaptive layer-wise growth harnesses quantum resources more effectively, avoiding barren plateaus and reducing measurement costs. In addition, although methods relying on continuous energy landscape optimization face greater challenges in this rugged landscape, relative to the continuous optimization method FEM, which claims generality for COPs without Ising reductions [shen2025free], sVQNHE offers superior global search in constraint-intensive, rugged landscapes. In contrast, FEM’s mean-field independence assumption ($\prod_{i}P_{i}(\sigma_{i})$) fails to capture the strong all-pair connectivity constraints inherent to cliques, leading to diffuse marginal distributions and higher post-processing losses [jagota1995approximating, shen2025free]. This underscores sVQNHE’s enhanced universality for combinatorial optimization problems.
These scaling experiments on MaxCut and Maximum Clique demonstrate that sVQNHE consistently delivers superior or competitive performance across diverse combinatorial optimization landscapes. On MaxCut, a problem with relatively smoother energy surfaces, sVQNHE achieves near-optimal solutions comparable to state-of-the-art classical solvers like FEM while offering slight advantages in rugged exploration via quantum correlations, together with substantial reductions in measurement overhead through NN-guided optimization. On Maximum Clique, a paradigmatically rugged, constraint-intensive problem with strong all-pair correlations, sVQNHE markedly outperforms all baselines and validates the critical value of its quantum module. Overall, sVQNHE emerges as a robust, scalable hybrid framework that leverages qubit-efficient encodings and synergistic quantum-classical computation to address both smooth and rugged combinatorial optimization problems effectively, positioning it as a promising candidate for near-term quantum utility in combinatorial optimization.